OpenStudy (anonymous):

Can someone please help me solve this equation? 7 + 3b = -2b/4

OpenStudy (anonymous):

Please, ASAP. :(

OpenStudy (uri):

options?

OpenStudy (anonymous):

a. b = -2 b. b = -1/2 c. b = 2 d. b = - 14/5

OpenStudy (anonymous):

-14/5*

OpenStudy (anonymous):

i will solve

OpenStudy (anonymous):

7=-3.5b

OpenStudy (anonymous):

b=7/-3,5=-2

OpenStudy (anonymous):

become a fan and give a medal

OpenStudy (anonymous):

Thank you all so much for helping me. I don't know what I'd do without you two. Maybe trying to solve this problem then getting it all wrong.. :/ but, thank you guys soooo much!!! Much appreciated. :)<3

OpenStudy (anonymous):

i am alone

OpenStudy (anonymous):

Thanks it was correct!!! <3

OpenStudy (uri):

@denis1206 o_O?

OpenStudy (anonymous):

for nothing sweety